Find the area enclosed by the curves f(x)=x^3 and g(x)=6x^2
Answer by Alan3354(69443)

Find the area enclosed by the curves f(x)=x^3 and g(x)=6x^2
------------
Find the intersections, if any.
---
x^3 - 6x^2 = 0
x = 0, x = 6
--> (0,0) and (6,216)
From 0 < x < 6 6x^2 is greater than x^3
--
INT(6x^2 - x^3) = 2x^3 - x^4/4 (Ignore the C)
Area = (2*6^3 - 6^4/4) - (0)
= 432 - 324
= 108 sq units